What is the geometric mean of 8 and 20
I did 8/x = x/20
xsquared = 8*20
x = sqrt160
x=12sqrt16
I don't know what I did wrong because the answer in the back of the book is not 12sqrt16 :( thank you for your help.

Answer by user_dude2008(1862)

8/x = x/20


x^2=160

x=sqrt(160)

x=sqrt(16*10)

x=sqrt(16)*sqrt(10)

x=4*sqrt(10)